Home‎ > ‎Crew 133‎ > ‎

112-GreenHabReport

XO Gartrelle to Cap Comm                        
The green hab is operating normally

The fan is operating normally at 68 degrees

The heater was turned down to 40 degrees F at 9AM
The heater was turned up to 75 degrees F at 5PM

All plants received water today. The herb seedlings still have not sprouted. The Mars shoot is still looking healthy. The lettuce sprouts have all emerged above ground and display leaves. They look very healthy. The pepper seedlings have not srouted as yet. A few are beginning to poke through the topsoil. 

The hydroponic seedlings were placed in the Greenhab at 9AM and returned to the upped hab at 5PM. A photo from this evening is attached. 

The run through of the hydroponic start up procedure was completed at 5:45 PM. Crewman Howell provided able assistance during this test and will participate in the actual hydroponic start-up. The test was successful and we have no major questions at this time. We will evaluate our water situation at midday tomorrow and decide if we have enough water to start hydroponic operations on Monday. Our next water re-supply is due Tuesday by 5PM. 

Greenhab readings for today are as follows:

                       Current         Min          Max
Temperature    77 Deg F    50 Deg F  86 Degrees F
Humidity          15%           19%         12%